Would you please help me with this problem ;0).
x over x-2, subtract x+1 over x = 8 over x^2-2x
I think the LCD is x^2-2x. I am lost from here. Thank you for your help.

Answer by stanbon(75887)

x over x-2, subtract x+1 over x = 8 over x^2-2x
--------------
x/(x-2) - (x+1)/x = 8/(x^2-2x)
Multiply thru by (x^2-2x) to get:
x(x) - (x+1)(x-2) = 8
x^2 - x^2+x+2 = 8
x+2 = 8
x = 6
------------
Checking:
6/4 - 7/6 = 8/24
36 - 28 = 8
8 - 8
=================
Cheers,
Stan H.